# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v2

@@ 3.3 Subsequential convergence
Let $\{(x^{k_{j}},y^{k_{j}},z^{k_{j}})\}$ be such that $(x^{k_{j}},y^{k_{j}},z^{k_{j}})\to(x^{*},y^{*},z^{*})$. From the proof of Proposition 3.15, we have that ${\left\|{y^{k}-x^{k-1}}\right\|}\to 0$ and ${\left\|{y^{k}-y^{k-1}}\right\|}\to 0$. It then follows that $y^{k_{j}}\to x^{*}$ by the triangle inequality, and therefore $x^{*}=y^{*}$. To show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, we first note that since $x^{*}=y^{*}$ and $z^{k}=x^{k}+\alpha\nabla f(x^{k})$ by (2.4), it follows that $z^{*}=y^{*}+\alpha\nabla f(y^{*})$. Since $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\leq\frac{1}{L_{{}_{\rm{f}}}}$, then $f+\frac{1}{2\alpha}\|\cdot\|^{2}$ is strongly convex by Remark 2.2 and therefore $y^{*}\in{\rm prox}_{\alpha f}(z^{*})$. Thus, $z^{*}=z^{*}+\tau(y^{*}-{\rm prox}_{\alpha f}(z^{*}))$. Hence, to show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, it remains to prove that $y^{*}\in P_{\Lambda}(y^{*},z^{*})$. To this end, note that by the convexity and continuity of $\frac{L_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p$ on $\operatorname*{dom}(g)$ (by Assumption 1.1(d)) and the formula (3.21), we have from (Beck, 2017, Theorem 3.16) and (3.21) that $\{\xi^{k_{j}}\}$ is bounded. Using again the continuity of $p$ on $\operatorname*{dom}(g)$ together with (2.2) and the fact that $\{y^{k}\}\subseteq\operatorname*{dom}(g)$, we see that accumulation points of $\{\xi^{k_{j}}\}$ belong to $\partial p(y^{*})$. Without loss of generality, we may assume that $\xi^{k_{j}}\to\xi^{*}$ where $\xi^{*}\in\partial p(y^{*})$.
@@
$\leq(f+h+p)(\bar{x})+\left\langle\nabla(f+h)(\bar{x})+\bar{\xi},x-\bar{x}\right\rangle+\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}}{2}{\left\|{x-\bar{x}}\right\|}^{2}+g(x),$
@@
$0\leq a+\left\langle v,x\right\rangle+\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}}{2}{\left\|{x}\right\|}^{2}+g(x)$ (3.64)
@@
for some $a\in{\rm I}\!{\rm R}$ and $v\in{\rm I}\!{\rm R}^{n}$. It then follows that $\liminf_{{\left\|{x}\right\|}\to\infty}\frac{g(x)}{{\left\|{x}\right\|}^{2}}\geq-\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}}{2}$ by dividing both sides of (3.64) by ${\left\|{x}\right\|}^{2}$ and using Cauchy-Schwarz inequality. By (Rockafellar and Wets, 1998, Exercise 1.24), the fact that $\frac{1}{\gamma}>L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\frac{1}{\beta}\geq L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}$ (that is, $\gamma<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+L_{{}_{\rm{p}}}}$), and (Rockafellar and Wets, 1998, Theorem 1.25), we know that $\{y^{k_{j}+1}\}$ given in (3.63) is bounded, with accumulation points lying in ${\rm prox}_{\gamma g}(u^{*})$. Since $y^{k_{j}+1}\to y^{*}$, it follows that $y^{*}\in{\rm prox}_{\gamma g}(u^{*})$, as desired.
To prove the last claim, we note first from (3.55) that $V_{k}\geq\Psi^{*}$. Summing (3.24) from 2 to $N+1$, we get
$\sum_{k=2}^{N+1}-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\sum_{k=2}^{N+1}\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}\leq V_{1}-\Psi^{*}.$ (3.65)
Meanwhile, we have $R(y^{k},z^{k})^{2}\leq{\left\|{(y^{k},z^{k})-(y^{k+1},z^{k+1})}\right\|}^{2}$ since $(y^{k+1},z^{k+1})\in T_{\Lambda}(y^{k},z^{k})$. These together with (3.24) and (3.56) give the desired result for the case $\beta<L_{{}_{\rm{p}}}^{-1}$. For the other case, the summation of (3.24) from 1 to $N$ gives
$\sum_{k=1}^{N}\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}{\left\|{y^{k}-y^{k+1}}\right\|}^{2}\leq V_{0}-\Psi^{*}.$
Using (3.65), we obtain
$\sum_{k=1}^{N}-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k+1}-x^{k}}\right\|}^{2}+\sum_{k=1}^{N}\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}{\left\|{y^{k}-y^{k+1}}\right\|}^{2}\leq V_{0}+V_{1}-2\Psi^{*},$
from where we can easily infer the result. ∎
3.4 Comments on the weak convexity assumption on $g$
We provide some remarks about the weak convexity assumption on $g$. First, this assumption is needed for guaranteeing the formula
$P_{\Lambda}(y,z)=\bigcup_{x\in{\rm prox}_{\alpha f}(z),\xi\in\partial p(y)}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,\Phi_{\Lambda,\xi}(w;y,z,x).$ (3.66)
in Lemma 3.2, but we can actually still ensure Eq. 3.66 without this weak convexity assumption on $g$ in some special cases, such as when either (I) $f=h=0$, or (II) $L_{{}_{\rm{p}}}=0$. In the former case, ${\rm prox}_{\alpha f}(z)\equiv z$ and $\gamma=\beta$, and consequently $P_{\Lambda}={\rm prox}_{\gamma g}(T_{\gamma p}(y))$, which is precisely the right-hand side of Eq. 3.66. On the other hand, when $L_{{}_{\rm{p}}}=0$, we have $\gamma=\alpha$ and hence
$P_{\Lambda}(y,z)$ $=\bigcup_{\begin{subarray}{c}x\in{\rm prox}_{\alpha f}(z)\\ \xi\in\partial p(y)\end{subarray}}{\rm prox}_{\gamma g}\left(x-\gamma\nabla f(x)-\gamma\nabla h(x)-\gamma\partial p(y)\right),$
@@ 3.4 Comments on the weak convexity assumption on $g$
$=\bigcup_{\begin{subarray}{c}x\in{\rm prox}_{\alpha f}(z)\\ \xi\in\partial p(y)\end{subarray}}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,g(w)+\frac{1}{2\gamma}{\left\|{w-x+\left(\gamma\nabla f(x)+\gamma\nabla h(x)+\gamma\xi\right)}\right\|}^{2},$
which is clearly equal to the right-hand side of (3.66).
We also note that when $g$ is not weakly convex, we can get a weaker version of Lemma 3.3. Indeed, from Eq. 3.7, we can obtain by setting $w=y$ that
$V_{\Lambda,\xi}(y,z,x)\leq Q_{\alpha(f+h)}(y;x)+p(y)+g(y)$ (3.67)
for all $(y,z)\in\operatorname*{dom}(g)\times{\rm I}\!{\rm R}^{n}$, $\xi\in\partial p(y)$ and $x\in{\rm prox}_{\alpha f}(z)$. The only difference is that we now lose the term $-\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}{\left\|{y-y^{+}}\right\|}^{2}$ in Lemma 3.3. Nevertheless, using this, we can obtain results parallel to Propositions 3.15 and 3.18 without requiring weak convexity, as long as the formula (3.66) holds true such as in cases (I) and (II) described above.
Proposition 3.19.
Suppose that Assumption 1.1(a), (b) and (d) are fulfilled, $g$ is a proper closed function, the formula (3.66) holds true, $\Psi$ has bounded level sets, and $\beta\in(0,L_{{}_{\rm{p}}}^{-1})$ if $L_{{}_{\rm{p}}}>0$. In addition, suppose that $L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}>0$, and $\alpha,\tau>0$ are chosen such that they satisfy (a), (b) or (c) in Proposition 3.15. If $\{(x^{k},y^{k},z^{k})\}$ is generated by Algorithm 1, then
  1. (i)
    $\{(x^{k},y^{k},z^{k})\}$ is bounded;
  2. (ii)
    ${\left\|{(x^{k},y^{k},z^{k})-(x^{k-1},y^{k-1},z^{k-1})}\right\|}\to 0$;
  3. (iii)
    If $(x^{*},y^{*},z^{*})$ is an accumulation point, then $x^{*}=y^{*}$ and $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$.
  4. (iv)
    There exists $\omega\in(0,\infty)$ (dependent on the stepsizes and the problem) such that the residual converges to 0 with rate
    $\min_{k=1,2,\dots,N}R(y^{k},z^{k})^{2}\leq\frac{\omega}{N}.$
@@
Using (3.67), Lemma 3.4, and (3.66), we can derive identical stepsize upper bounds for $\alpha$ as in Theorems 3.5, 3.9 and 3.12 to guarantee that the inequality
$V_{k-1}-V_{k}$$\geq\begin{cases}-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}&\text{if }L_{{}_{\rm{p}}}>0,\\ -\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}&\text{if }L_{{}_{\rm{p}}}=0,\end{cases}$ (3.68)
holds for all $k$, where $c(\alpha)$ is as defined in these theorems. With this, the proofs of parts (i) and (ii) are identical to that of Proposition 3.15. The proof of part (iii) is identical to the first part of the proof of Theorem 3.18, noting that weak convexity is not employed in the proof. For part (iv), when $L_{{}_{\rm{p}}}>0$, the result follows from the exact same arguments in Theorem 3.18 for this particular case, given the hypothesis that $\beta<L_{{}_{\rm{p}}}^{-1}$. Suppose now that $L_{{}_{\rm{p}}}=0$. From (3.68), we have for any $N\geq 0$ that
$\sum_{k=1}^{N+1}-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}\leq V_{0}-\Psi^{*}.$ (3.69)
On the other hand, from Eq. 3.3, we get $y^{k+1}-y^{k}=x^{k}-x^{k-1}+\tau^{-1}(z^{k+1}-z^{k})-\tau^{-1}(z^{k}-z^{k-1})$ It follows from the triangle inequality and the estimate $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ (for $a,b,c\in{\rm I}\!{\rm R}$) that
@@
We remark that in particular, our results apply to the exact same nonconvex setting of the DYS algorithm studied in (Bian and Zhang, 2021) (that is, without weak convexity assumption on $g$). Together with Remark 3.7 and Fig. 1, we have significantly improved the stepsize bounds derived by Bian and Zhang (2021) for the DYS algorithm.
Generalizing the above discussion, provided that the minimization of the function $\Phi_{\Lambda,\xi^{k}}(\cdot,y^{k},z^{k},x^{k})$ is not difficult or expensive to compute, we can simply replace the $y$-update rule (3.2) in Algorithm 1 with the formula
$y^{k+1}\in\bigcup_{x^{k}\in{\rm prox}_{\alpha f}(z^{k}),\xi^{k}\in\partial p(y^{k})}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\,\Phi_{\Lambda,\xi^{k}}(w;y^{k},z^{k},x^{k}).$ (3.70)
In this case, we can also apply the results to situations outside conditions (I) and (II) above. Similarly, we can obtain identical estimates for the stepsizes and derive results analogous to Proposition 3.19, where $T_{\Lambda}$ is as defined in (3.58) with $P_{\Lambda}$ given by the right-hand side of (3.66).
While finalizing this paper, we became aware of a recent preprint by Dao et al. (2024) that proposed to solve Eq. 1.1 with $L_{{}_{\rm{p}}}=0$ using an algorithm similar to ours. [^1] Specifically, setting $\theta=1$ in (Dao et al., 2024, Algorithm 1) coincides with our algorithm for the case $L_{{}_{\rm{p}}}=0$. Subsequential convergence was proved in (Dao et al., 2024) without the weak convexity assumption on $g$, which is also covered by our framework, in view of Proposition 3.19, as it corresponds to case (II) above. We highlight that another advantage of our analysis for this special case in which the settings coincide is that the upper bounds for stepsizes we computed are significantly larger than those in (Dao et al., 2024); see Figs. 2 and 3 for geometric illustrations. Meanwhile, (Dao et al., 2024) also proved global convergence of the full sequence under a Kurdyka-Łojasiewicz hypothesis, using the techniques by (Liu et al., 2019b), and a similar proof strategy can be employed to prove the same global convergence property for our algorithm.
@@ 4 Numerical experiments
In this section, we present experiments to corroborate the theoretical findings in the previous section. All experiments were conducted in MATLAB. For our method, we experiment with various values of $\tau$. For each $\tau$, we use the corresponding $\bar{\alpha}$ in Theorems 3.5, 3.9 and 3.12 times $9/10$ to ensure global subsequential convergence as suggested by Theorem 3.18. For comparing different methods that can fit in the framework of Algorithm 1, we report their required number of iterations and running time to reach $R(y,z)\leq 10^{-6}$.

### v4

@@ 3.3 Subsequential convergence
Let $\{(x^{k_{j}},y^{k_{j}},z^{k_{j}})\}$ be such that $(x^{k_{j}},y^{k_{j}},z^{k_{j}})\to(x^{*},y^{*},z^{*})$. From the proof of Proposition 3.15, we have that ${\left\|{y^{k}-x^{k-1}}\right\|}\to 0$ and ${\left\|{y^{k}-y^{k-1}}\right\|}\to 0$. It then follows that $y^{k_{j}}\to x^{*}$ by the triangle inequality, and therefore $x^{*}=y^{*}$. To show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, we first note that since $x^{*}=y^{*}$ and $z^{k}=x^{k}+\alpha\nabla f(x^{k})$ by (2.4), it follows that $z^{*}=y^{*}+\alpha\nabla f(y^{*})$. Since $\alpha<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}}\leq\frac{1}{L_{{}_{\rm{f}}}}$, then $f+\frac{1}{2\alpha}\|\cdot\|^{2}$ is strongly convex by Eq. 2.5 and therefore $y^{*}\in{\rm prox}_{\alpha f}(z^{*})$. Thus, $z^{*}=z^{*}+\tau(y^{*}-{\rm prox}_{\alpha f}(z^{*}))$. Hence, to show that $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$, it remains to prove that $y^{*}\in P_{\Lambda}(y^{*},z^{*})$. To this end, note that by the convexity and continuity of $\frac{\rho_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p$ on $\operatorname*{dom}(g)$ (by Assumption 1.1(d)) and the formula (3.13), we have from (Beck, 2017, Theorem 3.16) and (3.13) that $\{\xi^{k_{j}}\}$ is bounded. Using again the continuity of $p$ on $\operatorname*{dom}(g)$ together with (2.2) and the fact that $\{y^{k}\}\subseteq\operatorname*{dom}(g)$, we see that accumulation points of $\{\xi^{k_{j}}\}$ belong to $\partial p(y^{*})$. Without loss of generality, we may assume that $\xi^{k_{j}}\to\xi^{*}$ where $\xi^{*}\in\partial p(y^{*})$.
@@
$\leq(f+h+p)(\bar{x})+\left\langle\nabla(f+h)(\bar{x})+\bar{\xi},x-\bar{x}\right\rangle+\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\rho_{{}_{\rm{p}}}}{2}{\left\|{x-\bar{x}}\right\|}^{2}+g(x),$
@@
$0\leq a+\left\langle v,x\right\rangle+\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\rho_{{}_{\rm{p}}}}{2}{\left\|{x}\right\|}^{2}+g(x)$ (3.57)
@@
for some $a\in{\rm I}\!{\rm R}$ and $v\in{\rm I}\!{\rm R}^{n}$. It then follows that $\liminf_{{\left\|{x}\right\|}\to\infty}\frac{g(x)}{{\left\|{x}\right\|}^{2}}\geq-\frac{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\rho_{{}_{\rm{p}}}}{2}$ by dividing both sides of (3.57) by ${\left\|{x}\right\|}^{2}$ and using Cauchy-Schwarz inequality. By (Rockafellar and Wets, 1998, Exercise 1.24), the fact that $\frac{1}{\gamma}>L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\frac{1}{\beta}\geq L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\rho_{{}_{\rm{p}}}$ (that is, $\gamma<\frac{1}{L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}}+\rho_{{}_{\rm{p}}}}$), and (Rockafellar and Wets, 1998, Theorem 1.25), we know that $\{y^{k_{j}+1}\}$ given in (3.56) is bounded, with accumulation points lying in ${\rm prox}_{\gamma g}(u^{*})$. Since $y^{k_{j}+1}\to y^{*}$, it follows that $y^{*}\in{\rm prox}_{\gamma g}(u^{*})$, as desired. ∎
Remark 3.19.
It is not difficult to verify that when $\alpha<L_{{}_{\rm{f}}}^{-1}$, $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$ is equivalent to the conditions
@@
$\begin{cases}&y^{*}\in{\rm prox}_{\gamma g}\left(\frac{\gamma}{\alpha}\left(y^{*}-\alpha\nabla f(y^{*})-\alpha\nabla h(y^{*})\right)+\frac{\gamma}{\beta}T_{\beta p}(y^{*})\right),\\ &z^{*}=y^{*}+\alpha\nabla f(y^{*}).\end{cases}$ (3.58)
Meanwhile, in certain applications Lu and Zhou (2019); Lu et al. (2019), it is of interest to consider cases in which $p$ takes the form
$p(x)=-\max_{y\in\mathcal{Y}}\phi(x,y),$ (3.59)
where $\phi(\cdot,y)$ is continuously differentiable and convex for all $y\in\mathcal{Y}$, and $\mathcal{Y}$ is a compact set. This satisfies Assumption 1.1(d) with $\rho_{{}_{\rm{p}}}=0$. If $p$ takes this form and $f$, $h$, $g$ are convex and satisfy Assumption 1.1(a)-(c), then from (Lu et al., 2019, Proposition 1) and (3.58), it follows that any point $(y^{*},z^{*})\in\operatorname*{Fix}(T_{\Lambda})$ ensures that $y^{*}$ corresponds to a D-stationary point of Eq. 1.1, which, as discussed in Pang et al. (2017), is generally a stronger notion than the stationarity we used in Eq. 3.50. Together with Theorem 3.18, this shows that under these conditions on $f,g,h$, and $p$, our proposed algorithm is globally subsequentially convergent to a point $(y^{*},z^{*})$ where $y^{*}$ is a D-stationary point of (1.1).
The following theorem provides the $O(1/T)$ iteration complexity of Algorithm 1.
Theorem 3.20.
Suppose that the hypotheses of Proposition 3.15 hold. There exists $\omega>0$ such that $\min_{k=1,2,\dots,T}R(y^{k},z^{k})^{2}\leq\frac{\omega}{T},$ for all $T>1$.
@@
From (3.16), we have for any $N\geq 0$ that
$\sum_{k=1}^{N+1}-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}\leq V_{0}-\Psi^{*},$ (3.60)
where $\Psi^{*}\coloneqq\min_{x\in{\rm I}\!{\rm R}^{n}}\Psi(x)$. On the other hand, from Eq. 3.3, we get $y^{k+1}-y^{k}=x^{k}-x^{k-1}+\tau^{-1}(z^{k+1}-z^{k})-\tau^{-1}(z^{k}-z^{k-1}).$ It follows from the triangle inequality and the estimate $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ (for $a,b,c\in{\rm I}\!{\rm R}$) that
@@
While finalizing this paper, we became aware of a recent preprint by Dao et al. (2024) that proposed to solve Eq. 1.1 with $\rho_{{}_{\rm{p}}}=0$ using an algorithm similar to ours. [^6] Specifically, setting $\theta=1$ in (Dao et al., 2024, Algorithm 1) coincides with our algorithm for the case of $\rho_{{}_{\rm{p}}}=0$. The main advantage of our analysis for this coincided special case is that our upper bounds for stepsizes are significantly larger than those in (Dao et al., 2024); see Figs. 2 and 3 for geometric illustrations. Numerical results in Section 4.1.1 indicate that such a larger $\alpha$ is meaningful as it provides faster practical convergence. Meanwhile, subsequential convergence to stationary points is proven in both works, but ours allows for a wider range of stepsizes. (Dao et al., 2024) also proved global convergence of the full sequence under a Kurdyka-Łojasiewicz hypothesis, wherein they used the techniques by (Liu et al., 2019b). We note that a similar strategy can be employed to establish the same global convergence property for our algorithm with a wider range of stepsizes.
@@ 4 Numerical experiments
In this section, we present experiments to corroborate the theoretical findings in the previous section. All experiments were conducted in MATLAB. For our method, we experiment with various values of $\tau$. For each $\tau$, we use the corresponding $\bar{\alpha}$ in Theorems 3.5, 3.9 and 3.12, and we calculate $\alpha$ by multiplying $\bar{\alpha}$ by $0.99$ to ensure global subsequential convergence as suggested by Theorem 3.18. We also set $\beta=1/\rho_{{}_{\rm{p}}}$ according to our theoretical results when the $p$ term is present, and set it to $\infty$ otherwise as described in Section 3.1. The value of $\gamma$ is then decided by $\alpha$ and $\beta$ through Step 0 of Algorithm 1.
For a fair comparison of different methods that can fit in the framework of Algorithm 1, instead of Eq. 3.55, we report their respective required number of iterations and running time to reach an approximate first-order optimality measure independent of the algorithm. In particular, in view of (3.58), for any algorithm compared, regardless of their specific reformulation/transformation, we stop it when its main iterate $y^{k}$ satisfies
$\operatorname*{dist}\left(0,y^{k}-{\rm prox}_{\hat{\gamma}g}\left(\frac{\hat{\gamma}}{\hat{\alpha}}\left(y^{k}-\hat{\alpha}\nabla f(y^{k})-\hat{\alpha}\nabla h(y^{k})\right)+\frac{\hat{\gamma}}{\hat{\beta}}T_{\hat{\beta}p}(y^{k})\right)\right)\leq\epsilon$ (4.1)
for some $\epsilon>0$, where the choice of $f,g,h$, and $p$ are those used in our algorithm, and $(\hat{\alpha},\hat{\beta},\hat{\gamma})=(\alpha,\beta,\gamma)$ are calculated following the above procedure for our algorithm with $\tau=1$. From Remark 3.19, it is clear that at any stationary point $y^{*}$ of Eq. 1.1, the left-hand side of Eq. 4.1 vanishes, and thus it is a valid measure for the approximate first-order optimality measure. In all experiments in this section, we set $\epsilon=10^{-6}$.